# Legal Fact Prediction: The Missing Piece in Legal Judgment Prediction

### v2

@@ 4.3 Challenge & Bias Analysis
In Tables 2 and 3, we distinguish between cases where the defendant did or did not submit evidence. The results show that LFP-empowered LJP performs better in the former. This further suggests that counterevidence presented by the defendant can hinder LLMs’ ability to infer legal facts. Therefore, future research should focus on enhancing LLMs’ reasoning capabilities to better assess the authenticity of information.
@@
In Table 5, we select all cases that include both plaintiff and defendant evidence, sort the evidence list in different orders, and then feed them to the models for LFP-empowered LJP. We find that the order in which plaintiff and defendant evidence appears significantly influences the predicted judgments, introducing a bias favoring the party whose evidence is presented last. Specifically, when plaintiff evidence appears later, the LLMs are more likely to predict a complete or partial win for the plaintiff. Conversely, when defendant evidence appears later, the likelihood of the plaintiff losing the case increases. This bias may be attributed to the attention mechanism of LLMs Yu et al. (2024), which requires further exploration in future research.
@@
Figure 6 illustrates the impact of the difference in the number of evidence items between the plaintiff and the defendant on the predicted judgments. We observe that as the gap in evidence quantity increases, the LLMs generally tend to predict judgments with a higher complete-win rate and a lower partial-win rate compared to the ground truth. This suggests that an advantage in evidence quantity may lead LLMs to develop a bias toward fully supporting the plaintiff. However, the effect of this advantage on the loss rate is highly divergent: as the plaintiff’s advantage increases, the two closed-source models become less likely to predict a loss for the plaintiff, whereas the open-source models exhibit the opposite trend. Regardless of the model type, further efforts are needed to eliminate these biases and ensure that LLMs understand that an advantage in the number of evidence items does not necessarily indicate stronger evidence, nor does it have any causal relationship with the legal facts.
5 Discussion
Although the input for LFP is defined as the evidence list and the plaintiff’s claims, other trial-related information could also be incorporated into this task. As discussed by Medvedeva and Mcbride (2023), the ideal input for LJP should encompass any information available to the court or the parties at the time of performing LJP, such as complaints, defenses, and evidence submitted by the parties. This principle also applies to LFP. However, the information available to the court or the parties depends on the stage of the trial. For example, before filing a lawsuit, the plaintiff and the defendant may only have access to the evidence they personally possess. After filing, they gain access to each other’s evidence and arguments regarding the legal facts. In this work, we chose the evidence list as the basic input for LFP because, at different stages of the trial, both parties have access to certain evidence.
@@ 6 Related Work
Research on LJP can be traced back to the 1960s (Lawlor, 1963), which is one of the most fundamental tasks in legal AI. As judgment documents have become publicly accessible in many countries, researchers have extracted legal facts and judgment results from these documents, forming plenty of benchmark datasets for LJP research (e.g., (Xiao et al., 2018; Chalkidis et al., 2019; Malik et al., 2021; Semo et al., 2022; Chalkidis et al., 2022; Hwang et al., 2022)). Using judgment documents, numerous legal NLP studies have explored fact-based LJP, which predicts legal judgments based on legal facts, achieving promising predictive accuracy (e.g., (Luo et al., 2017; Zhong et al., 2018; Chen et al., 2019; Yue et al., 2021; Feng et al., 2022; Wu et al., 2022; Gan et al., 2023)). However, legal facts are not objective facts and are often difficult for LJP’s intended users to obtain before a judgment is rendered. Consequently, Medvedeva et al. (Medvedeva and Mcbride, 2023) recently pointed out that most existing LJP studies rely on unrealistic input such as legal facts, limiting their practicality. Several studies employed legal briefs (Tippett et al., 2021), complaint documents (McConnell et al., 2021), court debate records (Ma et al., 2021) for LJP, but these types of judicial documents are typically not publicly accessible, making it impractical to obtain large-scale datasets for training LJP models. To address the lack of practicality in current LJP research, this paper proposes the LFP task as a preliminary step to fact-based LJP. LFP-empowered LJP establishes a practical loop from evidence to judgment, making LJP more applicable in real-world scenarios.
@@
Legal Document Summarization (LDS) is the most correlated task with our LFP, which aims to automatically producing concise, accurate, and coherent summaries of legal texts (Kanapala et al., 2019), such as judgment documents Polsley et al. (2016), contracts Manor and Li (2019), and court debate records Duan et al. (2019). Intuitively, while LFP enriches and assembles fragmented and concise pieces of evidence into complete legal facts, LDS takes the opposite approach by refining legal texts and eliminating lengthy and complex details. Additionally, LFP requires inferring logically coherent factual information from conflicting or contradictory evidence, making it more challenging than text summarization.
@@ 7 Conclusion
This paper introduces the LFP task to automate the prediction of legal facts for the subsequent LJP task, addressing recent concerns that using ground-truth legal facts for LJP is impractical. We constructed a benchmark dataset for the LFP task, LFPBench, based on publicly available judicial documents. Extensive experiments conducted on LFPBench reveal that SOTA LLMs struggle to accurately determine legal facts when faced with conflicting or contradictory evidence. Moreover, these models exhibit biases related to the quantity and presentation order of evidence. Future work includes addressing the above limitations and constructing larger-scale LFP datasets to facilitate more extensive research.
@@
As the first step on the LFP task, this work has the following limitations. First, the evidence information in LFPBench is extracted from publicly accessible judgment documents in China, which are typically summarized and may lack many details of the original evidence, making it more challenging to predict legal facts. Second, the scope of case types covered by LFPBench remains limited, as it does not include criminal or administrative litigation cases.
@@
Our dataset is constructed based on judicial documents, which may present two ethical considerations: (1) Data Privacy and Confidentiality: Judicial documents often contain some basic personal information of litigants, such as names, addresses, and identity numbers. Although we have processed the data to remove or anonymize personally identifiable information (PII), we must still comply with data usage regulations and refrain from any de-anonymization attempts that could compromise personal privacy. (2) Bias and Fairness: Judicial datasets may inadvertently perpetuate biases present in historical legal decisions, related to gender, socioeconomic status, or other factors. Regular audits and updates can be conducted to minimize the risk of biased content.
Appendix A Additional Details on LFPBench
A.1 Annotation
@@ A.3 Copyright Issue
Appendix B Additional Details on Experiments
B.1 Documentation of LLMs
@@
B.2 Selection of Legal LLMs
@@
Table 8 reports the performance of the general-purpose LLM Llama3.1-Chinese-8B and five legal LLMs in the preliminary evaluation, including DISC-LawLLM Yue et al. (2024) and Lawyer-Llama-13B-V2 Huang et al. (2023), Law-Llama3.1-8B, LawJustice-Llama3.1-8B, and AIE-51-8-Law-Model lingminai (2025). The results show that while Llama3.1-Chinese-8B could fully follow our instructions, the long and complex instructions posed significant challenges for the legal LLMs. Finally, we selected the two legal LLMs with the highest success rate in instruction following for the main experiments in Section 4.
@@
B.3 Prompt Templates

### v3

@@ 4.3 Challenge & Bias Analysis (RQ2)
In Tables 4 and 6, we distinguish between cases where the defendant did or did not submit evidence. The results show that LFP-empowered LJP performs better in the former. This further suggests that counterevidence presented by the defendant can hinder LLMs’ ability to infer legal facts. Therefore, future research should focus on enhancing LLMs’ reasoning capabilities to better assess the authenticity of information.
@@
In Table 8, we select all cases that include both plaintiff and defendant evidence, sort the evidence list in different orders, and then feed them to the models for LFP-empowered LJP. We find that the order in which plaintiff and defendant evidence appears significantly influences the predicted judgments, introducing a bias favoring the party whose evidence is presented last. Specifically, when plaintiff evidence appears later, the LLMs are more likely to predict a complete or partial win for the plaintiff. Conversely, when defendant evidence appears later, the likelihood of the plaintiff losing the case increases. This bias may be attributed to the attention mechanism of LLMs yu2024mitigatepositionbiaslarge, which requires further exploration in future research.
@@
Figure 6 illustrates the impact of the difference in the number of evidence items between the plaintiff and the defendant on the predicted judgments. We observe that as the gap in evidence quantity increases, the LLMs generally tend to predict judgments with a higher complete-win rate and a lower partial-win rate compared to the ground truth. This suggests that an advantage in evidence quantity may lead LLMs to develop a bias toward fully supporting the plaintiff. However, the effect of this advantage on the loss rate is highly divergent: as the plaintiff’s advantage increases, the two closed-source models become less likely to predict a loss for the plaintiff, whereas the open-source models exhibit the opposite. Nevertheless, more efforts are needed to teach LLMs that more evidence items don’t necessarily mean stronger evidence or cause legal facts.
@@ 5 Related Work
Research on LJP can be traced back to the 1960s (lawlor1963computers), which is one of the most fundamental tasks in legal AI. As judgment documents have become publicly accessible in many countries, researchers have extracted legal facts and judgment results from these documents, forming plenty of benchmark datasets for LJP research (e.g., (xiao2018cail2018; chalkidis2019neural; malik2021ildc; semo2022classactionprediction; chalkidis2022lexglue; hwang2022multi)). Using judgment documents, numerous legal NLP studies have explored fact-based LJP, which predicts legal judgments based on legal facts, achieving promising predictive accuracy (e.g., (luo2017learning; zhong2018legal; chen2019charge; yue2021neurjudge; feng2022legal; wu2022towards; gan2023exploiting)). However, legal facts are not objective facts and are often difficult for LJP’s intended users to obtain before a judgment is rendered. Consequently, Medvedeva et al. (medvedeva2023legal) recently pointed out that most existing LJP studies rely on unrealistic input such as legal facts, limiting their practicality. Several studies employed legal briefs (tippett2021does), complaint documents (mcconnell2021case), court debate records (ma2021legal) for LJP, but these types of judicial documents are typically not publicly accessible, making it impractical to obtain large-scale datasets for training LJP models. To address the lack of practicality in current LJP research, this paper proposes the LFP task as a preliminary step to fact-based LJP. LFP-empowered LJP establishes a practical loop from evidence to judgment, thereby making LJP more applicable in real-world scenarios.
@@
Legal Document Summarization (LDS) is the most correlated task with our LFP, which aims to automatically producing concise, accurate, and coherent summaries of legal texts (kanapala2019text), such as judgment documents polsley2016casesummarizer, contracts manor2019plain, and court debate records duan2019legal. Intuitively, while LFP enriches and assembles fragmented and concise pieces of evidence into complete legal facts, LDS takes the opposite approach by refining legal texts and eliminating lengthy and complex details. Additionally, LFP requires inferring logically coherent factual information from conflicting or contradictory evidence, making it more challenging than text summarization.
@@ 6 Conclusion
This paper introduces the LFP task to automate the prediction of legal facts for the subsequent LJP task, addressing recent concerns that using ground-truth legal facts for LJP is impractical. We constructed a benchmark dataset for the LFP task, LFPBench, based on publicly available judicial documents. Extensive experiments conducted on LFPBench reveal that SOTA LLMs struggle to accurately determine legal facts when faced with conflicting or contradictory evidence, and exhibit biases related to the quantity and presentation order of evidence. Future work includes addressing the above limitations and constructing larger-scale LFP datasets to facilitate more extensive research.
@@
As the first step on the LFP task, this work has the following limitations. First, the evidence information in LFPBench is extracted from publicly accessible judgment documents in China, which are typically summarized and may lack some details of the original evidence, making it more challenging to predict legal facts. Second, the scope of case types covered by LFPBench remains limited, as it does not include criminal or administrative litigation cases. However, we would like to note that focusing on civil cases within a single jurisdiction is a common practice for benchmarking LJP (see Table 1).
@@
Our work may raise the following ethical considerations. (1) Data Privacy and Confidentiality: Judicial documents often contain some basic personal information of litigants, such as names, addresses, and identity numbers. Although we have processed the data to remove or anonymize personally identifiable information (PII), we must still comply with data usage regulations and refrain from any de-anonymization attempts that could compromise personal privacy. (2) Judicial Bias. Inappropriate applications of LJP may introduce ethical challenges, particularly because current fact-based LJP research often relies on legal facts extracted from court opinions, which may reflect judges or jurors’ biases. As a result, such biases can be embedded in LJP’s decisions. However, the introduction of the LFP task offers a way to alleviate this issue: it shifts the predictive foundation of LJP from biased legal facts in court opinions to facts predicted by LFP based on objective evidence. We believe this approach can, to some extent, reduce the influence of judicial bias. (3) Automated Adjudication. Some voices have proposed using LJP systems to replace human judges and juries, which has raised ethical concerns. However, we believe that the primary purpose of LFP and LJP is to assist litigants and their lawyers in predicting potential court opinions, thereby enabling them to adjust their strategies accordingly. This application can improve judicial transparency and reduce unnecessary judicial costs.
Additionally, we used ChatGPT to polish the writing and are responsible for all the materials presented in this work.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (62276077, 62376075, 62376076), JSPS Kakenhi (JP23K17456, JP23K25157, JP23K28096, JP25H01117, JP25K21207), and CREST (JPMJCR22M2).
Appendix A Discussion
Although the input for LFP is defined as the evidence list and the plaintiff’s claims, other trial-related information could also be incorporated into this task. As discussed by medvedeva2023legal, the ideal input for LJP should encompass any information available to the court or the parties at the time of performing LJP, such as complaints, defenses, and evidence submitted by the parties. This principle also applies to LFP. However, the information available to the court or the parties depends on the stage of the trial. For example, before filing a lawsuit, the plaintiff and the defendant may only have access to the evidence they personally possess. After filing, they gain access to each other’s evidence and arguments regarding the legal facts. In this work, we chose the evidence list as the basic input for LFP because, at different stages of the trial, both parties have access to certain evidence.
Appendix B Additional Details on LFPBench
B.1 Annotation
@@ B.3 Copyright Issue
Appendix C Additional Details on Experiments
C.1 Documentation of LLMs
@@
C.2 Selection of Legal LLMs
@@
Table 11 reports the performance of the general-purpose LLM Llama3.1-Chinese-8B and five legal LLMs in the preliminary evaluation, including DISC-LawLLM yue2024lawllm and Lawyer-Llama-13B-V2 Lawyer-LLama, Law-Llama3.1-8B, LawJustice-Llama3.1-8B, and AIE-51-8-Law-Model AIE. The results show that while Llama3.1-Chinese-8B could fully follow our instructions, the long and complex instructions posed significant challenges for the legal LLMs. Finally, we selected the two legal LLMs with the highest success rate in instruction following for the main experiments in Section 4.
@@
C.3 F1 Scores
We report the macro-average and micro-average F1 scores for various models and methods in Table 12. These results align with the accuracy metrics presented in Table 4.
Table 12: The macro-f1 and micro-f1 metrics of the ternary classification under different LJP approaches and models.
Model Evidence-based LFP-empowered Fact-based
Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1
GPT-4o 0.4957 0.5086 0.5073 0.5150 0.5503 0.5611
Claude3.5 0.4530 0.5083 0.4914 0.5264 0.5520 0.5780
Qwen2.5-14B 0.4109 0.4534 0.4335 0.4840 0.4725 0.5157
Llama3.1-Chinese-8B 0.3356 0.4101 0.3572 0.4125 0.3450 0.3971
Law-Llama3.1-8B 0.2825 0.3141 0.2859 0.3221 0.2523 0.2822
LawJustice-Llama3.1-8B 0.2714 0.3521 0.2289 0.2900 0.2422 0.3092
C.4 Prompt Templates